Calculate Kc for the following rxn:
2NO2 = N2 + 2O2; Kc= ?
Use the following data to find the unknown Kc:
1/2 N2 + 1/2 O2=NO ; Kc=4.8x10 -8
2NO2 = 2NO + O2 ; Kc =1.1x10 -3
1
Expert's answer
2017-10-29T11:53:07-0400
Kc = (p_(N_2 )×p_(O_2)^2)/(p_(〖NO〗_2)^2 ) – that’s what we have to calculate.
Kc1 = p_NO/(p_(N_2)^0.5×p_(O_2)^0.5 ) Kc2 = (p_NO^2×p_(O_2 ))/(p_(〖NO〗_2)^2 )
p_NO= Kc1 × p_(N_2)^0.5×p_(O_2)^0.5
Kc2 = (K_c1^2×p_(N_2 )×〖p_O^2〗_2)/(p_(〖NO〗_2)^2 ) = K_c1^2×K_c
Kc = K_c2/(K_c1^2 ) = ( 〖1.1×10〗^(-3))/(〖4.8×10〗^(-8) )^2 = ( 〖1.1×10〗^(-3))/〖23,04×10〗^(-16) = 4,7 × 1011
Answer
Kc = 4,7 × 1011